OpenStudy (hexagon001):

How to determine max|z exp(3 + iz^2)| for |z| ≤ 2

hartnn (hartnn):

|z| ≤ 2 ---> max |z| = 2 max|z exp(3 + iz^2)| = max|z exp(3) exp(iz^2)| = max|z exp(3 )|max |exp(iz^2)| now max value of e^(i theta) = 1 max|z exp(3 + iz^2)| = max|z exp(3 )|*1 = exp(3) max |z| = 2exp 3 got this ?

OpenStudy (hexagon001):

thanks